# noitemsep,topsep=0pt,parsep=0pt,partopsep=0pt

### v2

@@ 6.5 Algorithm Running Time
Figure 12 shows the running time for Alg. 1 (DistServe-Low) and Alg. 2 (DistServe-High) on a AWS m5d.metal instance with 96 cores as the number of GPUs ($N\times M$) provided to a single instance increases. According to the results, DistServe scales well with the number of GPUs and is independent of the model size. This is because the simulator only simulates discrete events and the running time is the same no matter how big the model is. On the other hand, both algorithms are highly parallelizable, as the searches for different parallelism strategies are independent of each other, allowing the execution time of the algorithms to accelerate almost linearly with more CPU cores.
@@ 7 Related Work
Inference serving. There has been plenty of work on inference serving recently. They range from general-purpose production-grade systems like TorchServe [6] and NVIDIA Triton [17] to systems optimized specifically for Transformer-based LLMs [29, 45, 8, 44, 19, 16, 51, 42]. Among them, Orca [45] introduces iteration-level scheduling to increase throughput. vLLM [28] proposes a novel memory management strategy for KVCache. SARATHI [8] suggests a chunked-prefill approach, splitting a prefill request into chunks and piggyback decoding requests to improve hardware utilization. FastServe [44] implements iteration-level preemptive scheduling to mitigate the queuing delay caused by long jobs. However, they all employ a colocation approach for prefill and decoding processing, thus leading to severe interference. There are also concurrent works Splitwise [33], TetriInfer [24], and DéjàVu [41] which adopt similar disaggregation idea to optimize LLM inference, further confirming the effectiveness of this method. Differently, DistServe emphasizes the goodput optimization scenario more and takes a closer look at the aspect of network bandwidth.
@@
Resource disaggregation. Resource disaggregated systems [38, 22, 15] decouple the hardware resources from the traditional monolithic server infrastructure and separate them into different pools to manage independently. It allows for more flexible, efficient, and scalable deployment and increases resource utilization. Many applications benefit from a truly disaggregated data center with high-speed network bandwidth and heterogenous hardware support  [48, 10]. DistServe adopts a similar concept by disaggregating its system components, allowing for independent resource scaling and management.
@@ 8 Conclusion
We present DistServe, a new LLM serving architecture that disaggregates the prefill and decoding computation. DistServe maximizes the per-gpu goodput – the maximum request rate that can be served adhering to the SLO attainment goal for each GPU provisioned, hence resulting in up to 4.48×4.48$\times$4.48 × lower cost per LLM query with guaranteed satisfaction of SLOs. Our findings affirm that as latency becomes an increasingly important metric for LLM services, prefill and decoding disaggregation is a vital strategy in promising improved performance and service quality guarantees.
@@ A.3 Decoding Phase Latency Modeling
Appendix B DistLLM Placements in End-to-end Experiments
In the end-to-end experiments 6.2, the tensor parallelism (TP) and pipeline parallelism (PP) configurations for prefill and decoding instances chosen by DistServe are listed on the right.

### v3

@@ 6.5 Algorithm Running Time
Figure 12 shows the running time for Alg. 1 (DistServe-Low) and Alg. 2 (DistServe-High) on an AWS m5d.metal instance with 96 cores as the number of GPUs ($N\times M$) provided to a single instance increases. According to the results, DistServe scales well with the number of GPUs and is independent of the model size. This is because the simulator only simulates discrete events and the running time is the same no matter how big the model is. On the other hand, both algorithms are highly parallelizable, as the searches for different parallelism strategies are independent of each other, allowing the execution time of the algorithms to accelerate almost linearly with more CPU cores.
@@
7 Discussion
In this paper, we focus on the goodput-optimized setting and propose DistServe under the large-scale LLM serving scenario. As LLMs are widely used and deployed across various service scenarios with different optimization targets and resource limits, it becomes almost impossible to find a one-size-fits-all solution that effectively addresses all aspects of LLM serving. In this section, we discuss the pros and cons of DistServe and potentially better solutions in other scenarios.
Throughput-optimized scenarios. In offline applications that are not latency-sensitive, users typically have lower requirements for response time [45]. This allows serving systems to shift focus towards maximizing overall throughput instead of goodput and the effectiveness of DistServe may be compromised. In this case, techniques such as chunked-prefill with piggyback [9, 3] may be preferred since it can fill each batch to the compute-bound threshold, thereby maintaining higher GPU utilization in every iteration.
Resource-constrained scenarios. Small-scale enterprises and individual researchers often lack the resources to deploy LLMs on large-scale clusters [45, 48]. In resource-constrained scenarios, such as environments with only a few or even a single GPU, the design space for DistServe is significantly limited. It struggles or even fails to adjust the parallel strategies and resource allocation to effectively enhance serving performance. In this case, simpler architectural choices like non-disaggregated systems [32, 3] may reduce deployment complexity and optimize operational efficiency.
Long-context scenarios. Nowadays, more and more GPT models support extremely long contexts, such as Claude-3 [11], Gemini-1.5 [22], and Large World Model (LWM) [34], which all have a 1M context window. In such scenarios, the transmission overhead will increase as the size of the KV cache grows linearly with the prompt length. However, the prefill computation grows quadratically, so the relative duration of transmission and prefill job decreases. Meanwhile, a longer context further exacerbates the disparity in computational demands between prefill and decoding jobs, leading to increased interference between them. Therefore, the disaggregation approach proposed in DistServe remains promising in long-context serving.
@@ 8 Related Work
Inference serving. There has been plenty of work on inference serving recently. They range from general-purpose production-grade systems like TorchServe [7] and NVIDIA Triton [19] to systems optimized specifically for Transformer-based LLMs [33, 54, 9, 53, 21, 18, 60, 50]. Among them, Orca [54] introduces continuous batching to increase throughput. vLLM [32] proposes paged-attention for fine-grained KV cache management. SARATHI [9] suggests a chunked-prefill approach, splitting a prefill request into chunks and piggybacking decoding requests to improve hardware utilization. FastServe [53] implements iteration-level preemptive scheduling to mitigate the queuing delay caused by long jobs. However, they all employ a colocation approach for prefill and decoding processing, thus leading to severe interference. There are also concurrent works such as Splitwise [38], TetriInfer [27] and DéjàVu [49] which adopt similar disaggregation idea to optimize LLM inference, further confirming the effectiveness of this method. Differently, DistServe emphasizes the goodput optimization scenario more and takes a closer look at the aspect of network bandwidth.
@@
Resource disaggregation. Resource disaggregated systems [43, 25, 17] decouple the hardware resources from the traditional monolithic server infrastructure and separate them into resource pools to manage independently. It allows for more flexible, efficient, and scalable deployment and increases resource utilization. Many applications benefit from a truly disaggregated data center with high-speed network bandwidth and heterogenous hardware support  [57, 12, 30]. DistServe shares the concept by disaggregating its system components, allowing for independent resource scaling and management.
@@ 9 Conclusion
We present DistServe, a new LLM serving architecture that disaggregates the prefill and decoding computation. DistServe maximizes the per-gpu goodput – the maximum request rate that can be served adhering to the SLO attainment goal for each GPU provisioned, hence resulting in up to 7.4×7.4$\times$7.4 × lower cost per LLM query with guaranteed satisfaction of SLOs. Our findings affirm that as latency becomes an increasingly important metric for LLM services, prefill and decoding disaggregation is a vital strategy in promising improved performance and service quality guarantees.
Acknowledgments. We sincerely thank our shepherd and the anonymous reviewers for their valuable feedback. This work was supported by the National Natural Science Foundation of China under the grant numbers 62172008, 62325201, and the National Natural Science Fund for the Excellent Young Scientists Fund Program (Overseas). Junda Chen is supported by UCSD fellowship and Hao Zhang is supported by UCSD faculty startup fund. Xin Jin is the corresponding author. Yinmin Zhong, Xuanzhe Liu, and Xin Jin are also with the Key Laboratory of High Confidence Software Technologies (Peking University), Ministry of Education.
@@ A.3 Decoding Phase Latency Modeling
(a) OPT-13B                                     (b) OPT-66B                                     (C) OPT-175B
(a) Code Completion                                             (b) Summarization
Appendix B DistServe Placements in End-to-end Experiments
Table 3 shows the tensor parallelism (TP) and pipeline parallelism (PP) configurations for prefill and decoding instances chosen by DistServe in the end-to-end experiments §6.2.
@@ Appendix B DistServe Placements in End-to-end Experiments
Table 3: The parallelism strategies chosen by DistServe in the end-to-end experiments.
Appendix C End-to-end Results under 99% SLO attainment
Figure 13 and Figure 14 show the end-to-end performance between DistServe and baselines with the same setup in §6.2 except that the SLO attainment goal is changed to 99%. We can see that under a more stringent SLO attainment goal, compared to vLLM, DistServe can still sustain 3×3$\times$3 ×8×8$\times$8 × higher rate and 1.24×1.24$\times$1.24 ×6.67×6.67$\times$6.67 × more stringent SLO. When compared to DeepSpeed-MII, DistServe can achieve 1.32×1.32$\times$1.32 ×8×8$\times$8 × higher rate and 1.20×1.20$\times$1.20 ×1.58×1.58$\times$1.58 × more stringent SLO.